You push the shopping cart uphill in the garage with the force Fp (icosα + jsinα) at ​​a distance of 15 m (icosθ + jsinθ), where the angle θ = 17 ∘ indicates the horizontal inclination of the floor. The absolute value of the thrust is Fp= 62.93 N and the direction α = -3.0 ∘ is slightly lowered and is also multiplied by the horizontal plane. As a result of pushing, the carts move at a steady speed. The weight of carts with shopping is 19 kg.

a)Determine how much work gravity Wmg will do for the cart?

b)Determine how much work WN the floor support will do for the cart?

c)Determine how much work WP you do on the cart
Answer #1

a) Force of gravity = mg (-j) = 19*9.81 (-j) = -186.39 (j)
Displacement = 15 m (icosθ + jsinθ) = 15 m (icos17 + jsin17)
Wmg = F.d = -186.39 (j) . 15 m (icos17 + jsin17) = -817.42 J

b) The normal force Fn is perpendicular to the direction of displacement and hence its dot product is zero. WN = 0 J

c) Force of thrust = Fp (icosα + jsinα) = 62.93 (icos(-3) + jsin(-3))
Displacement = 15 m (icosθ + jsinθ) = 15 m (icos17 + jsin17)
WP = F.d = 62.93 (icos(-3) + jsin(-3)) . 15 m (icos17 + jsin17) = 901.466-14.4439 = 887 J